it takes a florist 3 h 15 min to make 3 small centerpieces and 3 large centerpieces. it takes 6 h 20 min to make 4 small centerpieces and 7 large centerpieces. how long does it take to make each small centerpiece and each large centerpiece? write and solve a system of equations to find your answer.
Answer by ikleyn(52803)

.
Let x be the time for the florist to make 1 small centerpiece (in minutes), and
    
    y be the time for the florist to make 1 small centerpiece (in minutes).


Then from the condition you have this system of 2 equations in 2 unknowns.


    3x + 3y = 195   minutes   (1)    ( 105 minutes = 3 hours and 15 minutes )

    4x + 7y = 380   minutes   (2)    ( 380 minutes = 6 hours and 20 minutes )


Solve the system by the Elimination method in two steps.


STEP 1.   Divide the equation (1) (1) by 3 (both sides). Keep equation (2) as is. You will get the new system


     x +  y =  65             (3)   

    4x + 7y = 380             (4)      (eq4) is the same as eq(2) )  


STEP 2.  Multiply equation (1) by 4 (both sides). Then Subtract it from equation (4). You will get


     7y - 4y = 380 - 65*4

     3y      = 120           ==============>  y = 120/3 = 40 minutes.


Then from equation (1)  x = 65 - y = 65 - 40 = 25 minutes.


ANSWER.  25 minutes for small centerpiece  and  40 minutes for large centerpiece.

Solved.

---------------------

To see many other similar solved problems on systems of 2 equations in 2 unknowns see my lessons
    - Solution of the linear system of two equations in two unknowns by the Substitution method
    - Solution of the linear system of two equations in two unknowns by the Elimination method
    - Solution of the linear system of two equations in two unknowns using determinant
    - Geometric interpretation of the linear system of two equations in two unknowns
    - Useful tricks when solving systems of 2 equations in 2 unknowns by the Substitution method
    - Solving word problems using linear systems of two equations in two unknowns

    - Oranges and grapefruits
    - Using systems of equations to solve problems on tickets
    - Three methods for solving standard (typical) problems on tickets
    - Using systems of equations to solve problems on shares
    - Using systems of equations to solve problems on investment
    - Two mechanics work on a car
    - The Robinson family and the Sanders family each used their sprinklers last summer
    - Roses and vilolets
    - Counting calories and grams of fat in combined food
    - A theater group made appearances in two cities
    - HOW TO algebreze and solve this problem on 2 equations in 2 unknowns
    - One unusual problem to solve using system of two equations
    - Solving mentally word problems on two equations in two unknowns

    - OVERVIEW of lessons on solving systems of two linear equations in two unknowns

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.